# Eigen Attention: Attention in Low-Rank Space for KV Cache Compression

### v1

@@ Abstract
Utkarsh Saxena Purdue University saxenau@purdue.edu                        Gobinda Saha Purdue University gsaha@purdue.edu                        Sakshi Choudhary Purdue University choudh23@purdue.edu                        Kaushik Roy Purdue University kaushik@purdue.edu
@@ 1 Introduction
The recent boom in artificial intelligence applications and their widespread public adoption can be attributed to the human-like capabilities of large language models (LLMs). LLMs have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks lmsys.org (2024). The maximum number of tokens these models can process simultaneously to understand and generate text is referred to as the context/sequence length, and this closely determines their performance limits. Hence, there has been considerable interest in increasing the context length to enhance their capabilities Zhang et al. (2024); Ding et al. (2024); Achiam et al. (2023). Longer context lengths open up new possibilities, such as summarizing lengthy documents, retrieving information to answer questions about extensive texts, and analyzing code. To make applications enabled by LLMs more accessible, developing techniques to serve these models efficiently is crucial. One standard technique to accelerate LLM inference on GPUs is caching the intermediate attention keys and values through a KV cache to avoid expensive re-computation for every generated token. However, it is observed that at long context lengths, the KV cache becomes the new memory and latency bottleneck Pope et al. (2022). Furthermore, while batching multiple requests together amortizes the weight access cost of LLMs, it exacerbates the KV cache memory overhead. Consider the weight and KV cache memory footprint for the Llama-2 7B model with a batch size of 16 and a context length of 32k tokens. Here, the weight memory occupies 14 GB, while the KV cache requires a significantly higher memory of 256 GB at 16-bit precision. In essence, increasing the context/sequence length of LLMs has transformed LLM inference into a memory-bound problem. The entire KV cache must be bought on-chip from the off-chip GPU memory for each newly generated token while the computation core stays idle, waiting for data.
@@ 3 Related Works
KV Cache Compression. As mentioned in Section 2, the KV cache size is computed as $2*b*n*d*h*L*p$. KV cache compression methods target these factors to reduce its overall memory footprint. Multi-query attention Shazeer (2019) and grouped query attention Ainslie et al. (2023) reduce the number of attention heads $h$. Quantization based methods reduce the precision $p$ Yang et al. (2024); Kang et al. (2024); Zirui Liu et al. (2024); Hooper et al. (2024). Notably, works like KIVI Zirui Liu et al. (2024) and KV Quant Hooper et al. (2024) have demonstrated that the precision of $\mathbf{K}$ and $\mathbf{V}$ matrices can be reduced to as low as 2-bits. Several works attempt to reduce the sequence length $n$ by only caching $\mathbf{K}$ and $\mathbf{V}$ corresponding to a subset of tokens Beltagy et al. (2020). Another strategy is to cache K and V according to token importance Zhang et al. (2023); Adnan et al. (2024); Liu et al. (2024); Niu et al. (2024). $\textnormal{H}_{2}\textnormal{O}$ Zhang et al. (2023) finds important tokens by monitoring accumulated attention scores. KeyFormer Adnan et al. (2024) improves $\textnormal{H}_{2}\textnormal{O}$ by considering the importance of discarded tokens as well. Recently, Wu and Tu (2024) demonstrated that the cached $\mathbf{K}$ and $\mathbf{V}$ can be reused across the decoder layers, essentially reducing $L$. In contrast to these techniques, Eigen Attention reduces the embedding dimension $d$ of each cached $\mathbf{K}$ and $\mathbf{V}$ vector. It is orthogonal to the existing KV cache compression techniques and can be used in conjunction with them.
@@ 4.3 Rotational Position Embedding
Model Parameters KV Cache PPL $\downarrow$ Acc $\uparrow$ Wikitext C4 PiQA WinoGrande Arc-e Arc-c HellaSwag Avg-Acc OPT 30b Baseline 9.56 10.69 0.78 0.68 0.70 0.35 0.54 0.61 0.8x 9.61 (+0.05) 10.75 (+0.06) 0.78 0.67 0.70 0.34 0.54 0.61 (-0.00) 0.7x 9.94 (+0.38) 10.98 (+0.29) 0.78 0.66 0.69 0.35 0.54 0.60 (-0.01) 0.6x 10.80 (+1.24) 11.47 (+0.78) 0.76 0.64 0.69 0.32 0.52 0.59 (-0.02) 66b Baseline 9.34 10.28 0.79 0.69 0.72 0.37 0.56 0.63 0.8x 9.36 (+0.02) 10.29 (+0.01) 0.79 0.69 0.72 0.37 0.56 0.63 (-0.00) 0.7x 9.51 (+0.17) 10.40 (+0.12) 0.78 0.68 0.72 0.37 0.56 0.62 (-0.01) 0.6x 9.68 (+0.34) 10.54 (+0.26) 0.78 0.68 0.70 0.36 0.55 0.62 (-0.01) MPT 7b Baseline 7.68 9.60 0.79 0.69 0.75 0.41 0.57 0.64 0.8x 8.02 (+0.34) 10.14 (+0.54) 0.79 0.68 0.74 0.39 0.56 0.63 (-0.01) 0.7x 8.45 (+0.77) 10.80 (+1.20) 0.78 0.68 0.72 0.38 0.54 0.62 (-0.02) 0.6x 9.61 (+1.93) 12.26 (+2.66) 0.76 0.68 0.69 0.35 0.52 0.60 (-0.04) 30b Baseline 6.40 8.44 0.80 0.70 0.79 0.47 0.60 0.67 0.8x 6.48 (+0.08) 8.55 (+0.11) 0.81 0.71 0.79 0.47 0.60 0.68 (+0.01) 0.7x 6.66 (+0.26) 8.82 (+0.38) 0.80 0.71 0.78 0.46 0.59 0.67 (-0.00) 0.6x 7.01 (+0.61) 9.38 (+0.94) 0.79 0.70 0.77 0.44 0.57 0.66 (-0.01) Llama-2 7b Baseline 5.47 6.97 0.78 0.69 0.76 0.43 0.57 0.65 0.8x 5.96 (+0.49) 7.82 (+0.85) 0.76 0.66 0.72 0.40 0.54 0.61 (-0.04) 0.7x 6.28 (+0.81) 8.55 (+1.58) 0.75 0.63 0.70 0.38 0.52 0.60 (-0.05) 0.6x 7.48 (+2.01) 10.07 (+3.10) 0.74 0.63 0.65 0.33 0.48 0.56 (-0.09) 13b Baseline 4.88 6.47 0.79 0.72 0.79 0.48 0.60 0.68 0.8x 5.06 (+0.18) 6.77 (+0.30) 0.78 0.72 0.78 0.45 0.59 0.66 (-0.02) 0.7x 5.32 (+0.44) 7.19 (+0.72) 0.77 0.70 0.77 0.45 0.57 0.65 (-0.03) 0.6x 6.10 (+1.22) 9.34 (+2.87) 0.73 0.65 0.70 0.36 0.48 0.58 (-0.10) 70b Baseline 3.32 5.52 0.82 0.78 0.83 0.54 0.65 0.72 0.8x 3.44 (+0.12) 5.74 (+0.22) 0.81 0.76 0.81 0.51 0.63 0.71 (-0.01) 0.7x 3.60 (+0.28) 5.95 (+0.43) 0.81 0.76 0.81 0.51 0.61 0.70 (-0.02) 0.6x 3.78 (+0.46) 6.23 (+0.71) 0.80 0.75 0.81 0.50 0.59 0.69 (-0.03) Llama-3 8b Baseline 6.14 8.88 0.80 0.73 0.80 0.51 0.60 0.69 0.8x 6.72 (+0.58) 9.93 (+1.05) 0.78 0.73 0.78 0.48 0.57 0.67 (-0.02) 0.7x 7.22 (+1.08) 10.91 (+2.03) 0.78 0.71 0.77 0.45 0.55 0.65 (-0.04) 0.6x 8.51 (+2.37) 13.20 (+4.32) 0.76 0.67 0.74 0.42 0.50 0.62 (-0.07) 70b Baseline 2.85 6.73 0.82 0.81 0.87 0.60 0.66 0.75 0.8x 3.22 (+0.37) 7.12 (+0.39) 0.82 0.80 0.85 0.59 0.66 0.74 (-0.01) 0.7x 3.59 (+0.74) 7.48 (+0.75) 0.82 0.79 0.84 0.57 0.65 0.73 (-0.02) 0.6x 5.54 (+2.69) 10.39 (+3.66) 0.78 0.73 0.77 0.45 0.54 0.65 (-0.10)
@@ 5.1 Setup
We evaluate Eigen Attention across three model families: OPT Zhang et al. (2022), MPT MosaicML-MPT , and Llama Touvron et al. (2023); Llama- (3), each with distinct position encoding schemes. OPT employs learnable position embeddings, MPT utilizes AliBi Press et al. (2021), and the Llama model family employs RoPE Su et al. (2024). We conduct evaluations on both language generation and zero-shot tasks. The language generation tasks include perplexity evaluation on Wikitext-2 Merity et al. (2016) and C4 Dodge et al. (2021) datasets. The zero-shot tasks are obtained from lm-eval-harness framework Gao et al. (2023): PiQA Bisk et al. (2020), Winogrande Sakaguchi et al. (2021), Arc-easy/challenge Clark et al. (2018), and HellaSwag Zellers et al. (2019).
@@
Model KV Cache PPL $\downarrow$ Acc $\uparrow$ Wikitext C4 PiQA WinoGrande Arc-e Arc-c HellaSwag Avg-Acc MPT-7b Baseline 7.68 9.6 0.79 0.69 0.75 0.41 0.57 0.64 0.6x 9.21 (+1.53) 11.44 (+2.14) 0.79 0.69 0.72 0.39 0.55 0.63 (-0.01) Llama-2-7b Baseline 5.47 6.97 0.78 0.69 0.76 0.43 0.57 0.65 0.6x 6.55 (+1.07) 8.55 (+1.58) 0.78 0.67 0.74 0.41 0.54 0.63 (-0.02) Llama-2-13b Baseline 4.88 6.47 0.79 0.72 0.79 0.48 0.60 0.68 0.6x 5.64 (+0.76) 7.77 (+1.30) 0.78 0.68 0.75 0.44 0.58 0.65 (-0.03) Llama-3-8b Baseline 6.14 8.88 0.80 0.73 0.80 0.51 0.60 0.69 0.6x 7.60 (+1.47) 11.44 (+2.56) 0.80 0.70 0.79 0.48 0.58 0.67 (-0.02)
@@ 5.2 Results
Model Baseline Eigen Attn 0.6x OPT-66b 1.57 ms 0.62 ms (-60%) Llama-2-70b 0.79 ms 0.77 ms (-3%) Llama-3-70b 0.74 ms 0.80 ms (+8%)
Figure 3: Perplexity on Wikitext with different KV cache sizes in GB ($n$ = 2048) obtained via different quantization precision and group size. For Eigen Attention, we compress the KV cache to 0.6x and then apply quantization.
@@
Latency Comparisons: We evaluate the latency of the attention operation during the token generation phase for generating 128 tokens with a 2048-token synthetic prompt on 2 NVIDIA A100 (80 GB) GPUs, as shown in Table 4. We observe an impressive 60% latency improvement for the OPT-66b model, which can be attributed to the FLOPs reduction from Eigen Attention and the reduced latency in fetching the compressed KV cache from memory. For models with RoPE embedding (i.e., Llama-2 and Llama-3), we perform an additional transformation before the attention dot product (Figure 6), which diminishes the latency gains achieved with Eigen Attention. For Llama-2-70b, we observe only a 3% latency improvement, while for Llama-3-70b, there is a latency penalty of 8%.

### v2

@@ Abstract
Utkarsh Saxena,    Gobinda Saha,    Sakshi Choudhary,    Kaushik Roy Purdue University {saxenau, gsaha, choudh23, kaushik}@purdue.edu
@@ 1 Introduction
The recent boom in artificial intelligence applications and their widespread public adoption can be attributed to the human-like capabilities of large language models (LLMs). LLMs have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks lmsys.org (2024). The maximum number of tokens these models can process simultaneously to understand and generate text is referred to as the context/sequence length, and this closely determines their performance limits. Hence, there has been considerable interest in increasing the context length to enhance their capabilities Zhang et al. (2024a); Ding et al. (2024); Achiam et al. (2023). Longer context lengths open up new possibilities, such as summarizing lengthy documents, retrieving information to answer questions about extensive texts, and analyzing code. To make applications enabled by LLMs more accessible, developing techniques to serve these models efficiently is crucial. One standard technique to accelerate LLM inference on GPUs is caching the intermediate attention keys and values through a KV cache to avoid expensive re-computation for every generated token. However, it is observed that at long context lengths, the KV cache becomes the new memory and latency bottleneck Pope et al. (2022). Furthermore, while batching multiple requests together amortizes the weight access cost of LLMs, it exacerbates the KV cache memory overhead. Consider the weight and KV cache memory footprint for the Llama-2 7B model with a batch size of 16 and a context length of 32k tokens. Here, the weight memory occupies 14 GB, while the KV cache requires a significantly higher memory of 256 GB at 16-bit precision. In essence, increasing the context/sequence length of LLMs has transformed LLM inference into a memory-bound problem. The entire KV cache must be bought on-chip from the off-chip GPU memory for each newly generated token while the computation core stays idle, waiting for data.
@@ 3 Related Works
KV Cache Compression. As mentioned in Section 2, the KV cache size is computed as $2*b*n*d*h*L*p$. KV cache compression methods target these factors to reduce its overall memory footprint. Multi-query attention Shazeer (2019) and grouped query attention Ainslie et al. (2023) reduce the number of attention heads $h$. Quantization based methods reduce the precision $p$ Yang et al. (2024); Kang et al. (2024); Zirui Liu et al. (2024); Hooper et al. (2024). Notably, works like KIVI Zirui Liu et al. (2024) and KV Quant Hooper et al. (2024) have demonstrated that the precision of $\mathbf{K}$ and $\mathbf{V}$ matrices can be reduced to as low as 2-bits. Several works attempt to reduce the sequence length $n$ by only caching $\mathbf{K}$ and $\mathbf{V}$ corresponding to a subset of tokens Beltagy et al. (2020). Another strategy is to cache K and V according to token importance Zhang et al. (2023); Adnan et al. (2024); Liu et al. (2024); Niu et al. (2024); Li et al. (2024); Zhang et al. (2024b). $\textnormal{H}_{2}\textnormal{O}$ Zhang et al. (2023) finds important tokens by monitoring accumulated attention scores. KeyFormer Adnan et al. (2024) improves $\textnormal{H}_{2}\textnormal{O}$ by considering the importance of discarded tokens as well. Recently, Wu and Tu (2024) demonstrated that the cached $\mathbf{K}$ and $\mathbf{V}$ can be reused across the decoder layers, essentially reducing $L$. In contrast to these techniques, Eigen Attention reduces the embedding dimension $d$ of each cached $\mathbf{K}$ and $\mathbf{V}$ vector. It is orthogonal to the existing KV cache compression techniques and can be used in conjunction with them.
@@ 4.3 Rotational Position Embedding
Model Params KV Cache PPL $\downarrow$ Acc $\uparrow$ Wikitext C4 PiQA WinoG Arc-e Arc-c HellaS Avg-Acc OPT 30b Baseline 9.56 10.69 0.78 0.68 0.70 0.35 0.54 0.61 0.8x 9.61 (+0.05) 10.75 (+0.06) 0.78 0.67 0.70 0.34 0.54 0.61 (-0.00) 0.7x 9.94 (+0.38) 10.98 (+0.29) 0.78 0.66 0.69 0.35 0.54 0.60 (-0.01) 0.6x 10.80 (+1.24) 11.47 (+0.78) 0.76 0.64 0.69 0.32 0.52 0.59 (-0.02) 66b Baseline 9.34 10.28 0.79 0.69 0.72 0.37 0.56 0.63 0.8x 9.36 (+0.02) 10.29 (+0.01) 0.79 0.69 0.72 0.37 0.56 0.63 (-0.00) 0.7x 9.51 (+0.17) 10.40 (+0.12) 0.78 0.68 0.72 0.37 0.56 0.62 (-0.01) 0.6x 9.68 (+0.34) 10.54 (+0.26) 0.78 0.68 0.70 0.36 0.55 0.62 (-0.01) MPT 7b Baseline 7.68 9.60 0.79 0.69 0.75 0.41 0.57 0.64 0.8x 8.02 (+0.34) 10.14 (+0.54) 0.79 0.68 0.74 0.39 0.56 0.63 (-0.01) 0.7x 8.45 (+0.77) 10.80 (+1.20) 0.78 0.68 0.72 0.38 0.54 0.62 (-0.02) 0.6x 9.61 (+1.93) 12.26 (+2.66) 0.76 0.68 0.69 0.35 0.52 0.60 (-0.04) 30b Baseline 6.40 8.44 0.80 0.70 0.79 0.47 0.60 0.67 0.8x 6.48 (+0.08) 8.55 (+0.11) 0.81 0.71 0.79 0.47 0.60 0.68 (+0.01) 0.7x 6.66 (+0.26) 8.82 (+0.38) 0.80 0.71 0.78 0.46 0.59 0.67 (-0.00) 0.6x 7.01 (+0.61) 9.38 (+0.94) 0.79 0.70 0.77 0.44 0.57 0.66 (-0.01) Llama-2 7b Baseline 5.47 6.97 0.78 0.69 0.76 0.43 0.57 0.65 0.8x 5.96 (+0.49) 7.82 (+0.85) 0.76 0.66 0.72 0.40 0.54 0.61 (-0.04) 0.7x 6.28 (+0.81) 8.55 (+1.58) 0.75 0.63 0.70 0.38 0.52 0.60 (-0.05) 0.6x 7.48 (+2.01) 10.07 (+3.10) 0.74 0.63 0.65 0.33 0.48 0.56 (-0.09) 13b Baseline 4.88 6.47 0.79 0.72 0.79 0.48 0.60 0.68 0.8x 5.06 (+0.18) 6.77 (+0.30) 0.78 0.72 0.78 0.45 0.59 0.66 (-0.02) 0.7x 5.32 (+0.44) 7.19 (+0.72) 0.77 0.70 0.77 0.45 0.57 0.65 (-0.03) 0.6x 6.10 (+1.22) 9.34 (+2.87) 0.73 0.65 0.70 0.36 0.48 0.58 (-0.10) 70b Baseline 3.32 5.52 0.82 0.78 0.83 0.54 0.65 0.72 0.8x 3.44 (+0.12) 5.74 (+0.22) 0.81 0.76 0.81 0.51 0.63 0.71 (-0.01) 0.7x 3.60 (+0.28) 5.95 (+0.43) 0.81 0.76 0.81 0.51 0.61 0.70 (-0.02) 0.6x 3.78 (+0.46) 6.23 (+0.71) 0.80 0.75 0.81 0.50 0.59 0.69 (-0.03) Llama-3 8b Baseline 6.14 8.88 0.80 0.73 0.80 0.51 0.60 0.69 0.8x 6.72 (+0.58) 9.93 (+1.05) 0.78 0.73 0.78 0.48 0.57 0.67 (-0.02) 0.7x 7.22 (+1.08) 10.91 (+2.03) 0.78 0.71 0.77 0.45 0.55 0.65 (-0.04) 0.6x 8.51 (+2.37) 13.20 (+4.32) 0.76 0.67 0.74 0.42 0.50 0.62 (-0.07) 70b Baseline 2.85 6.73 0.82 0.81 0.87 0.60 0.66 0.75 0.8x 3.22 (+0.37) 7.12 (+0.39) 0.82 0.80 0.85 0.59 0.66 0.74 (-0.01) 0.7x 3.59 (+0.74) 7.48 (+0.75) 0.82 0.79 0.84 0.57 0.65 0.73 (-0.02) 0.6x 5.54 (+2.69) 10.39 (+3.66) 0.78 0.73 0.77 0.45 0.54 0.65 (-0.10)
@@ 5.1 Setup
We evaluate Eigen Attention across three model families: OPT Zhang et al. (2022), MPT MosaicML-MPT , and Llama Touvron et al. (2023); Llama- (3), each with distinct position encoding schemes. OPT employs learnable position embeddings, MPT utilizes AliBi Press et al. (2021), and the Llama model family employs RoPE Su et al. (2024). We conduct evaluations on both language generation and zero-shot tasks. The language generation tasks include perplexity evaluation on Wikitext-2 Merity et al. (2016) and C4 Dodge et al. (2021) datasets. The zero-shot tasks are obtained from lm-eval-harness framework Gao et al. (2023): PiQA Bisk et al. (2020), Winogrande (WinoG) Sakaguchi et al. (2021), Arc-easy/challenge Clark et al. (2018), and HellaSwag (HellaS) Zellers et al. (2019).
@@
Model KV Cache PPL $\downarrow$ Acc $\uparrow$ Wikitext C4 PiQA WinoG Arc-e Arc-c HellaS Avg-Acc MPT-7b Baseline 7.68 9.6 0.79 0.69 0.75 0.41 0.57 0.64 0.6x 9.21 (+1.53) 11.44 (+2.14) 0.79 0.69 0.72 0.39 0.55 0.63 (-0.01) Llama-2-7b Baseline 5.47 6.97 0.78 0.69 0.76 0.43 0.57 0.65 0.6x 6.55 (+1.07) 8.55 (+1.58) 0.78 0.67 0.74 0.41 0.54 0.63 (-0.02) Llama-2-13b Baseline 4.88 6.47 0.79 0.72 0.79 0.48 0.60 0.68 0.6x 5.64 (+0.76) 7.77 (+1.30) 0.78 0.68 0.75 0.44 0.58 0.65 (-0.03) Llama-3-8b Baseline 6.14 8.88 0.80 0.73 0.80 0.51 0.60 0.69 0.6x 7.60 (+1.47) 11.44 (+2.56) 0.80 0.70 0.79 0.48 0.58 0.67 (-0.02)
Figure 3: PPL on Wikitext with different KV cache sizes in GB ($n$ = 2048) obtained via different quantization precision and group size. For Eigen Attention, we compress the KV cache to 0.6x and then apply quantization.
@@ 5.2 Results
Table 4: Comparison of Eigen Attention with H2O. Combination of Eigen Attention and H2O leads to the best tradeoff between accuracy and KV cache size.
Model Method KV Cache Avg. 0-shot Acc $\uparrow$ Llama-2-7b H2O 0.2x 0.54 H2O 0.4x 0.64 Eigen Attn 0.6x 0.63 Eigen Attn + H2O 0.24x 0.62 Llama-3-8b H2O 0.2x 0.56 H2O 0.4x 0.67 Eigen Attn 0.6x 0.67 Eigen Attn + H2O 0.24x 0.65
@@
Model Baseline Attn Eigen Attn 0.6x OPT-66b 1.57 ms 0.62 ms (-60%) Llama-2-70b 0.79 ms 0.77 ms (-3%) Llama-3-70b 0.74 ms 0.80 ms (+8%)
@@
Eigen Attention with token pruning: We compare Eigen Attention with a popular token pruning approach H2O Zhang et al. (2023), which only keeps K and V corresponding to a subset of tokens within the KV Cache. Table 4 shows the average zero shot accuracy comparison at different KV cache compression ratios. We observe that H2O at 0.4x KV cache achieves similar accuracy to Eigen Attention at 0.6x KV cache. Superior performance of H2O over Eigen Attention can be attributed to the dynamic nature of the algorithm. While Eigen Attention compresses KV cache based on statistics derived offline, H2O is able to achieve better generalization by making decisions to eject tokens from KV cache at runtime. However, both Eigen Attention and H2O compress KV cache along different dimensions and can be combined to achieve further compression as seen in Table 4. The combination of Eigen Attention and H2O achieves the best performance highlighting the orthogonality of both approaches. By reducing the hidden dimension of K and V by 60% using Eigen Attention and using H2O to keep only 40% of tokens in KV cache, we are able to compress KV cache to 0.24x while achieving accuracy comparable to Eigen Attention at 0.6x and H2O at 0.4x KV cache.
Table 6: Total inference latency (in seconds) for Llama-3-8b model at different prompt and generation length.
Batch Size Prompt tokens Generated tokens Baseline Attn Eigen Attn (0.6x) 1 8192 2048 76 82 (+8%) 1 8192 16384 785 880 (+12%) 1 8192 32768 2048 2350 (+13%) 32 2048 512 180 156 (-13%) 32 2048 2048 826 859 (+4%) 32 2048 4096 OOM 2366
Latency Comparisons: We compare the latency of Eigen Attention and standard attention baseline for different models. In Table 5, we show average latency per attention layer during the token generation phase for generating 128 tokens with a 2048-token synthetic prompt on 2 NVIDIA A100 GPUs. We observe an impressive 60% latency improvement for the OPT-66b model, which can be attributed to the FLOPs reduction from Eigen Attention and the reduced latency in fetching the compressed KV cache from memory. For models with RoPE embedding (i.e., Llama-2-70b and Llama-3), we perform an additional transformation before the attention dot product (Figure 6), which diminishes the latency gains achieved with Eigen Attention. For Llama-2-70b, we observe only a 3% latency improvement, while for Llama-3-70b, there is a latency penalty of 8%. We further analyze end to end inference latency at different context lengths and batch sizes in Table 6 for Llama-3-8b on a single NVIDIA A100 GPU. We see that increasing the context length increases the latency overhead with Eigen Attention due to the presence of additional transformation to manage RoPE embedding. For the maximum sequence length of 40k, there is a 13% latency overhead with Eigen Attention for Llama-3-8b. When batch size is increased, the KV cache size increases which increases the memory bounded nature of inference. In this scenario, additional transformation required by Eigen Attention leads to much lesser latency overhead and often faster inference than baseline attention. Most notably, at high batch size and context length, baseline attention leads to out of memory (OOM) error on GPU while Eigen Attention does not. This highlights the main contribution of our work: to enable long context inference by compressing KV cache.
Table 7: Perplexity and zero-shot accuracy when using different calibration datasets for MPT-7b model.
Calib. dataset KV Cache Wikitext (PPL) C4 (PPL) Zero-shot Acc. (%) C4 0.8 8.03 9.96 0.63 0.7 8.57 10.47 0.63 0.6 9.63 11.44 0.62 Alpaca 0.8 8.12 10.07 0.64 0.7 3.02 10.88 0.62 0.6 10.86 12.42 0.61